OpenStudy (anonymous):

Jason and Kyle both choose a number from 1 to 10 at random. What is the probability that both numbers are odd?

OpenStudy (anonymous):

5 out of 10 numbers are odd: 1, 3, 5, 7, 9:\[\frac{ 5 }{ 10 } * \frac{ 5 }{ 10 }= \frac{ 1 }{ 4 }\]